# Privacy-Preserving Resilient Vector Consensus

### v1

@@ 4.2 Distribution of Final Value
From Theorem 1, although the expectation of the final value is in the convex hull of the normal agents’ initial states, we cannot guarantee that each single convergence result always lies in the convex hull. This uncertainty arises because the specific distribution of the final value remains unknown. Thus, we investigate the accuracy of each final value in terms of its residuals, i.e., the deviation from the expected value. Specifically, we propose to use the Mahalanobis distance [25] and the multivariate Chebyshev’s inequality [26] to analyze the residuals. The details are provided as follows.
@@
$D_{M}(z)=\sqrt{(z-\mathbb{E}(z))^{T}\Sigma^{-1}(z-\mathbb{E}(z))}.$ (11)
In resilient vector consensus, our primary concern is the convergence accuracy of the random vector $x(\infty)$. The commonly used Euclidean distance is not suitable in this scenario due to varying variances of each component in $x(\infty)$ and existing correlations among them. These factors introduce different measurement scales across dimensions. Therefore, we opt for the Mahalanobis distance to quantify the deviation of the final value from its expected value. This metric offers a comprehensive approach to distance measurement by incorporating variances and correlations between components. By mitigating the influence of disparate scales on distance calculation, the Mahalanobis distance provides a more precise evaluation.
@@
We then show that covariance matrix $x(\infty)$ is non-singular. Combining Definition 4.4 and Lemma 4.5, we can derive the result below.
@@
Under PP-ADRC, if equation (6) holds, $\overline{\mathcal{G}}(0),\overline{\mathcal{G}}(1),\overline{\mathcal{G}}(2),\ldots$ is repeatedly reachable, and the covariance matrix of $x(\infty)$ is non-singular, then
$\mathrm{P}\left\{[D_{M}(x(\infty))]^{2}\leq\chi\right\}\geq 1-\frac{d}{\chi},\quad\forall\chi>0.$
@@
$\left.[\Sigma]_{kk^{\prime}}=\left\{\begin{array}[]{ll}\operatorname{var}(x_{i,k}(\infty))&\text{if }k=k^{\prime}\\ \tau_{k,k^{\prime}}\sqrt{\operatorname{var}(x_{i,k}(\infty))\operatorname{var}(x_{i,k^{\prime}}(\infty))}&\text{if }k\neq k^{\prime},\end{array}\right.\right.$ (13)
where $\tau_{k,k^{\prime}}$ is the correlation coefficient. Consequently, we can derive the determinant of $\Sigma$, i.e., $|\Sigma|=h(\tau_{1,2},\ldots,\tau_{d-1,d})\operatorname{var}(x_{i,1}(\infty))\times\cdots\times\operatorname{var}(x_{i,d}(\infty))$, where $h(\cdot)$ represents a polynomial. It means that the volume is directly proportional to the variances.
For the case where the covariance matrix of $x(\infty)$ is singular, we can obtain the following results [28].
@@
For a random vector $z\in\mathbb{R}^{d}$, if its covariance matrix $\Sigma$ is singular, then there exists at least one non-zero row vector $\zeta$ such that $\operatorname{var}(\zeta z)=0$, which means $\zeta z$ is a constant.
@@
Lemma 4.8 implies that certain components of the random variable can be expressed as linear polynomials of the remaining components. Identifying these extraneous components allows us to obtain a random vector $Z^{\prime}\in\mathbb{R}^{d^{\prime}}$ with a non-singular covariance matrix $\Sigma^{\prime}$. We consider the case where the covariance matrix of $x(\infty)$ is singular and without loss of generality, the latter $d-d^{\prime}$ components of $x(\infty)$ are supposed to be extraneous. It means that Theorem 5.20 still holds for the remaining $d^{\prime}$ components, where the result simply degenerates into a $d^{\prime}$-dimensional ellipsoid (or a line if $d^{\prime}=1$).
@@ 4.3 Convex Hull Change
In contrast to scalar consensus, the high-dimensional convex hull is defined by numerous vertices, dimensions, and coefficients. Moreover, the stochastic nature of noise can significantly alter the convex hull of typical agents, making it challenging to characterize. Furthermore, following a single iteration, the new convex hull may exhibit no overlap with the original configuration.
To solve this issue, we consider a special case that only partial dimensions need to be protected. Consequently, the convex hull of dimensions without adding noise retains its original form. Our focus shifts to analyzing the transformation of the convex hull in the remaining dimensions of the state. It is worth noting that analyzing high-dimensional convex hulls can be challenging, while the $1$-dimensional final value $x_{i,k}(\infty)$ can be assessed by examining its expectation and variance. Hence, we start convex hull change analysis from the perspective of one dimension.
We first evaluate the expectation and variance of $x_{i,k}(\infty)$. Then, we determine the one-dimensional convergence accuracy using Chebyshev’s inequality. Finally, we quantize the change of the convex hull caused by the Gaussian noise and its associated probability. The properties of the random variable $x_{i,k}(\infty)$ are given below.
@@
Under PP-ADRC, if equation (6) holds, and $\overline{\mathcal{G}}(0),\overline{\mathcal{G}}(1),\overline{\mathcal{G}}(2),\ldots$ is repeatedly reachable, then
$\mathbb{E}\left[x_{i,k}(\infty)\right]=\psi[\overline{\mathbf{x}}_{0}]^{k},\operatorname{var}(x_{i,k}(\infty))\leq\frac{\lambda^{2}}{1-\upsilon^{2}},1\leq k\leq d,$
@@
Proof 4.11.
Similar to that in [20], using the independence and zero-mean of noise, we can obtain
$\mathbb{E}\left[x_{i,k}(\infty)\right]=\lim\limits_{t\rightarrow\infty}\mathbb{E}\left[\Psi_{t+1,0}[\overline{\mathbf{x}}_{0}]^{k}\right]=\psi[\overline{\mathbf{x}}_{0}]^{k}.$
@@
$\operatorname{var}(x_{i,k}(\infty))$ $\leq\frac{1}{\overline{n}}\lim\limits_{t\rightarrow\infty}\sum_{q=0}^{t}\sum_{j=1}^{\overline{n}}\left[\Psi_{t+1,q}^{T}\Psi_{t+1,q}\right]_{jj}\text{var}([\mathbf{v}]_{jk}(q))$
@@
Under PP-ADRC, if equation (6) holds, and $\overline{\mathcal{G}}(0),\overline{\mathcal{G}}(1),\overline{\mathcal{G}}(2),\ldots$ is repeatedly reachable, then
$\mathrm{P}\{|x_{i,k}(\infty)-\psi[\overline{\mathbf{x}}_{0}]^{k}|\leq(l_{k}+r_{k})\}$ $\geq 1-\frac{\operatorname{var}(x_{i,k}(\infty))}{(l_{k}+r_{k})^{2}}$ (14)
@@
$\mathrm{P}\{|x_{i,k}(\infty)-\psi[\overline{\mathbf{x}}_{0}]^{k}|\leq(l_{k}+r_{k})\}$ $\geq 1-\frac{\operatorname{var}(x_{i,k}(\infty))}{(l_{k}+r_{k})^{2}}$
@@
Given that $\psi$ is a row vector with elements summing to one, it follows that $\psi[\overline{\mathbf{x}}_{0}]^{k}$ must lie within the range bounded by the maximum and minimum values of $[\overline{\mathbf{x}}_{0}]^{k}$. Consequently, Lemma 4.12 implies that $x_{i,k}(\infty)$ lies within the interval $[\overline{\mathbf{x}}_{0}]^{k}_{\operatorname{max}}+r_{k}$ to $[\overline{\mathbf{x}}_{0}]^{k}_{\operatorname{min}}-r_{k}$ with a probability of at least $1-\frac{\lambda^{2}/(1-\upsilon^{2})}{(l_{k}+r_{k})^{2}}$.
@@
where $10<\beta<$ 1, $r_{k}$ is a freely chosen parameter for $k=1,2,\ldots,\beta d$, and $\mathcal{A}$ is the convex hull of the initial states of the normal agents. Under PP-ADRC, if equation (6) holds, $\overline{\mathcal{G}}(0),\overline{\mathcal{G}}(1),\overline{\mathcal{G}}(2),\ldots$ is repeatedly reachable, and $(1-\beta)$ fraction of the dimensions of the state vector are not added with noises, the probability $\mathrm{P}\{x(\infty)\in\mathcal{C}\}$ that the final value lies within the convex hull $\mathcal{C}$ satisfies
$\mathrm{P}\{x(\infty)\in\mathcal{C}\}\geq\prod\limits_{k=1}^{\beta d}\left[1-\frac{\lambda^{2}/(1-\upsilon^{2})}{(l_{k}+r_{k})^{2}}\right].$ (17)
Proof 4.18.
@@
First, we assume $b_{i}$ is a constant and then find the lower bound of $\mathrm{dist}(b,c)$. Apparently when $c_{i}=b_{i}$, $\mathrm{dist}(b,c)$ is lower bounded, which is $\min\limits_{c_{i}}(\mathrm{dist}(b,c))=\sqrt{\sum\limits_{i=1}^{\kappa}\left[b_{i}^{2}(r_{1}^{2}+\ldots+r_{\beta d}^{2})\right]}$. Then, the problem becomes finding the upper bound of $\min\limits_{c_{i}}(\mathrm{dist}(b,c))$, which is denoted by $\max\limits_{b_{i}}\min\limits_{c_{i}}(\mathrm{dist}(b,c))$. Given that $b_{1}+\ldots+b_{\kappa}=1$ and $b_{1}^{2}+\ldots+b_{\kappa}^{2}\leq(b_{1}+\ldots+b_{\kappa})^{2}$, we can know that the upper bound of $b_{1}^{2}+\ldots+b_{\kappa}^{2}$ is $1$ when $b_{i_{0}}=1,\ i_{0}\in{1,2,\ldots,\kappa}\ \text{ and }\ b_{i}=0,\ i\neq i_{0}$. Therefore, we can obtain $\max\limits_{b_{i}}\min\limits_{c_{i}}(\mathrm{dist}(b,c))=\sqrt{r_{1}^{2}+\ldots+r_{\beta d}^{2}}$. According to the definition, $\max\limits_{b_{i}}\min\limits_{c_{i}}(\mathrm{dist}(b,c))$ is exactly the Hausdorff distance $D_{H}(\mathcal{B},\mathcal{C})$ between $\mathcal{B}$ and $\mathcal{C}$. So we can obtain
@@
$\mathrm{P}\{x(\infty)\in\mathcal{C}\}$ $\geq\prod\limits_{k=1}^{\beta d}\left[1-\frac{\operatorname{var}(x_{i,k}(\infty)}{(l_{k}+r_{k})^{2}}\right]$
@@ 5 Privacy Analysis
In this section, we analyze the $\rho-$concentrated geo-privacy of PP-ADRC and compare it with $(\varepsilon,\delta)-$differential privacy. At last, we make a detailed discussion on the relationship between convergence accuracy and privacy.
5.1 $\rho-$Concentrated Geo-Privacy
We first investigate the $\rho-$CGP properties of PP-ADRC. As defined in Definition 8, in the resilient vector consensus, the input is the initial states of normal agents $\overline{\mathbf{x}}_{0}$ and the output is the infinite sequence of noised states $\mathbf{Y}=\{\overline{\mathbf{y}}(t)\}_{t=0}^{\infty}$.
@@ 5.1 $\rho-$Concentrated Geo-Privacy
Under PP-ADRC, if equation (6) holds, and $\overline{\mathcal{G}}(0),\overline{\mathcal{G}}(1),\overline{\mathcal{G}}(2),\ldots$ is repeatedly reachable, then we achieve $\rho-$CGP, where $\rho=\bigg{[}\frac{n\upsilon^{2}}{2\lambda^{2}[\upsilon^{2}-(1-\gamma_{l})^{2}]}\bigg{]}$.
Proof 5.21.
@@
Then, we need to show that based on the above definitions, the outputs $\mathbf{Y}_{k,\overline{\mathbf{x}}_{0}}(\mathbf{N}_{k})$ and $\mathbf{Y}_{k,\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N}_{k}^{\prime})$ have the same distribution. For the normal agents, we use the method of mathematical induction. From equation (3), we can derive $y_{i}^{\prime}(0)=x_{i}^{\prime}(0)+\eta_{i}^{\prime}(0)=x_{i}(0)+\delta_{i}+\eta_{i}(0)-\delta_{i}=y_{i}(0),i=1,2,\cdots,\overline{n}$. We can conclude that for every normal agent $i$, the calculated centerpoints $s_{i}(0)$ and $s_{i}^{\prime}(0)$ are the same. According to equation (4), we can derive $x_{i}^{\prime}(1)=x_{i}(1)+(1-\gamma_{i}(0))\delta_{i}$. By induction, we can easily obtain that $x_{i}^{\prime}(h)=x_{i}(h)+\prod_{t=0}^{h-1}\left[1-\gamma_{i}(t)\right]\delta_{i},h=1,2,\cdots,k$, followed by $y_{i}(h)=y_{i}^{\prime}(h)$. It means that $\mathbf{Y}_{k,\overline{\mathbf{x}}_{0}}(\mathbf{N}_{k})=\mathbf{Y}_{k,\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N}_{k}^{\prime})$, indicating that $\mathbf{N}_{k}^{\prime}\in R_{k}^{\prime}$ and we successfully build a bijective correspondence. For any $\mathbf{N}_{k}^{\prime}\in R_{k}^{\prime}$, there exists $(\mathbf{N}_{k},\Delta\mathbf{N}_{k})\in R_{k}\times(\mathbb{R}^{\overline{n}\times d})^{k+1}$ such that $\mathbf{N}_{k}^{\prime}=\mathbf{N}_{k}+\Delta\mathbf{N}_{k}$. Therefore, we derive the integral by changing variables $\mathrm{P}\{\mathbf{N}\in\Omega|\mathbf{Y}_{\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N})\in\mathcal{O}\}=\lim\limits_{k\to\infty}\int_{R_{k}}\mathsf{f}_{\overline{n}(k+1)}(\mathbf{N}_{k}+\Delta\mathbf{N}_{k})\mathrm{d}\mathbf{N}_{k},$ indicating that the distribution $\mathbf{y}_{\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N})$ can be expressed as $\lim\limits_{k\rightarrow\infty}\mathsf{f}_{\overline{n}(k+1)}(\mathbf{N}_{k}+\Delta\mathbf{N}_{k})$.
@@
$=\lim_{k\to\infty}\frac{1}{\alpha-1}\mathrm{log}\bigg{[}\int_{R_{k}}\prod\limits_{h=0}^{k}\prod\limits_{i=1}^{\overline{n}}\bigg{[}\frac{1}{\sqrt{(2\pi\sigma^{2}(h))^{d}}}\mathrm{exp}\left(-\frac{||\eta_{i}(h)-(1-\alpha)\Delta\eta_{i}(h)||^{2}_{2}-||(1-\alpha)\Delta\eta_{i}(h)||^{2}_{2}+(1-\alpha)||\Delta\eta_{i}(h)||^{2}_{2}}{2\sigma^{2}(h)}\right)\bigg{]}$
$\mathrm{d}\eta_{1}(h)\cdots\eta_{\overline{n}}(h)$
$=\lim_{k\to\infty}\frac{1}{\alpha-1}\mathrm{log}\prod\limits_{h=0}^{k}\{\prod\limits_{i=1}^{\overline{n}}\underset{\eta_{i}(h){\sim}\mathcal{MVN}((1-\alpha)\Delta\eta_{i}(h),\sigma^{2}(h)I_{d})}{\mathbb{E}}\bigg{[}\mathrm{exp}\left(-\frac{-||(1-\alpha)\Delta\eta_{i}(h)||^{2}_{2}+(1-\alpha)||\Delta\eta_{i}(h)||^{2}_{2}}{2\sigma^{2}(h)}\right)\bigg{]}\}$
$=\lim_{k\to\infty}\sum\limits_{h=0}^{k}\bigg{[}\sum\limits_{i=1}^{\overline{n}}\frac{\alpha||\Delta\eta_{i}(h)||^{2}_{2}}{2\lambda^{2}\upsilon^{2h}}\bigg{]}\leq\bigg{[}\frac{\alpha n}{2\lambda^{2}}+\lim_{k\to\infty}\sum\limits_{h=1}^{k}\frac{\alpha n(1-\gamma_{l})^{2h}}{2\lambda^{2}\upsilon^{2h}}\bigg{]}\mathrm{dist}(\overline{\mathbf{x}}_{0},\overline{\mathbf{x}}_{0}^{\prime})^{2}$
$\leq\alpha\bigg{[}\frac{n\upsilon^{2}}{2\lambda^{2}[\upsilon^{2}-(1-\gamma_{l})^{2}]}\bigg{]}\mathrm{dist}(\overline{\mathbf{x}}_{0},\overline{\mathbf{x}}_{0}^{\prime})^{2},$
where $\mathcal{MVN}((1-\alpha)\Delta\eta_{i}(h),\sigma^{2}(h)I_{d})$ represents a multivariate Gaussian distribution with mean $(1-\alpha)\Delta\eta_{i}(h)$ and covariance matrix $\sigma^{2}(h)I_{d}$. Consequently, we have found the upper bound of the Rényi divergence, thereby completing the proof.

### v2

@@ 4.2 Distribution of Final Value
From Theorem 4.2, although the expectation of the final value is in the convex hull of the normal agents’ initial states, we cannot guarantee that each single convergence result always lies in the convex hull. This uncertainty arises because the specific distribution of the final value remains unknown. Thus, we investigate the accuracy of each final value in terms of its residuals, i.e., the deviation from the expected value. Specifically, we propose to use the Mahalanobis distance [38] and the multivariate Chebyshev’s inequality [39] to analyze the residuals. The details are provided as follows.
@@
$D_{M}(z)=\sqrt{(z-\mathbb{E}(z))^{\top}\Sigma^{-1}(z-\mathbb{E}(z))}.$ (13)
In resilient vector consensus, our primary concern is the convergence accuracy of the random vector $\xi$. The commonly used Euclidean distance is not suitable in this scenario due to varying variances of each component in $\xi$ and existing correlations among them. These factors introduce different measurement scales across dimensions. Therefore, we opt for the Mahalanobis distance to quantify the deviation of the final value from its expected value. This metric offers a comprehensive approach to distance measurement by incorporating variances and correlations between components. By mitigating the influence of disparate scales on distance calculation, the Mahalanobis distance provides a more precise evaluation.
Here, we provide the details of the multivariate Chebyshev’s inequality [39].
@@
If the covariance matrix of $\xi$ is non-singular, then by Definition 4.6 and Lemma 4.7, we obtain the following result.
@@
Under PP-ADRC, if the condition in (6) holds, $\overline{\mathcal{G}}(0),\overline{\mathcal{G}}(1),\overline{\mathcal{G}}(2),\ldots$ is repeatedly reachable, and the covariance matrix of $\xi$ is non-singular, then
$\mathrm{P}\left\{[D_{M}(\xi)]^{2}\leq\chi\right\}\geq 1-\frac{d}{\chi},\quad\forall\chi>0.$
@@
$\left.[\Sigma]_{kk^{\prime}}=\left\{\begin{array}[]{ll}\operatorname{var}(\xi_{k})&\text{if }k=k^{\prime}\\ \tau_{k,k^{\prime}}\sqrt{\operatorname{var}(\xi_{k})\operatorname{var}(\xi_{k^{\prime}})}&\text{if }k\neq k^{\prime},\end{array}\right.\right.$ (15)
where $\tau_{k,k^{\prime}}$ is the correlation coefficient. Consequently, we can derive the determinant of $\Sigma$, i.e., $|\Sigma|=h(\tau_{1,2},\ldots,\tau_{d-1,d})\operatorname{var}(\xi_{1})\times\cdots\times\operatorname{var}(\xi_{d})$, where $h(\cdot)$ represents a polynomial. It means that the volume is directly proportional to the variances.
For the case where the covariance matrix of $\xi$ is singular, we can obtain the following result [41].
@@
For a random vector $z\in\mathbb{R}^{d}$, if its covariance matrix $\Sigma$ is singular, then there exists at least one non-zero row vector $\zeta$ such that $\operatorname{var}(\zeta z)=0$, implying $\zeta z$ is a constant.
@@
Lemma 4.10 implies that certain components of the random variable can be expressed as linear polynomials of the remaining components. Identifying these extraneous components allows us to obtain a random vector $Z^{\prime}\in\mathbb{R}^{d^{\prime}}$ with a non-singular covariance matrix $\Sigma^{\prime}$. We consider the case where the covariance matrix of $\xi$ is singular and without loss of generality, the latter $d-d^{\prime}$ components of $\xi$ are supposed to be extraneous. It means that Theorem 4.8 still holds for the remaining $d^{\prime}$ components, where the result simply degenerates into a $d^{\prime}$-dimensional ellipsoid (or a line if $d^{\prime}=1$).
@@ 4.3 Convex Hull Change
In contrast to scalar consensus, the high-dimensional convex hull is defined by numerous vertices, dimensions, and coefficients. Moreover, the stochastic nature of noise can significantly alter the convex hull of the normal agents, making it challenging to characterize. Furthermore, following a single iteration, the new convex hull may exhibit no overlap with the original configuration.
To solve this issue, we consider a special case that only partial dimensions need to be protected. Consequently, the convex hull of dimensions without adding noise retains its original form. Our focus shifts to analyzing the transformation of the convex hull in the remaining dimensions of the state. It is worth noting that analyzing high-dimensional convex hulls can be challenging, while the $1$-dimensional final value $\xi_{k}$ can be assessed by examining its expectation and variance. Hence, we start the convex hull change analysis from the perspective of one dimension.
We first evaluate the expectation and variance of $\xi_{k}$. Then, we determine the one-dimensional convergence accuracy using Chebyshev’s inequality. Finally, we quantize the change of the convex hull caused by Gaussian noise and its associated probability. The properties of $\xi_{k}$ are given below.
@@
Under PP-ADRC, if the condition in (6) holds, and $\overline{\mathcal{G}}(0),\overline{\mathcal{G}}(1),\overline{\mathcal{G}}(2),\ldots$ is repeatedly reachable, then
$\mathbb{E}\left[\xi_{k}\right]=\psi[\overline{\mathbf{x}}_{0}]^{k},\operatorname{var}(\xi_{k})\leq\frac{\lambda^{2}}{1-\upsilon^{2}},1\leq k\leq d,$
@@
Proof 4.13.
Following [24], by using the independence and zero mean of the noise, along with the strong ergodicity (indicating that the limit of $\Psi_{t+1,0}$ exists), we can obtain that
$\mathbb{E}\left[\xi_{k}\right]=\lim\limits_{t\rightarrow\infty}\mathbb{E}\left[[\Psi_{t+1,0}]_{i}[\overline{\mathbf{x}}_{0}]^{k}\right]=\psi[\overline{\mathbf{x}}_{0}]^{k}.$
@@
$\operatorname{var}(\xi_{k})$ $\leq\frac{1}{\overline{n}}\lim\limits_{t\rightarrow\infty}\sum_{q=0}^{t}\sum_{j=1}^{\overline{n}}\left[\Psi_{t+1,q}^{T}\Psi_{t+1,q}\right]_{jj}\text{var}([\mathbf{v}]_{jk}(q))$
@@
Under PP-ADRC, if the condition in (6) holds, and $\overline{\mathcal{G}}(0),\overline{\mathcal{G}}(1),\overline{\mathcal{G}}(2),\ldots$ is repeatedly reachable, then
$\mathrm{P}\{|\xi_{k}-\psi[\overline{\mathbf{x}}_{0}]^{k}|\leq(l_{k}+r_{k})\}$ $\geq 1-\frac{\operatorname{var}(\xi_{k})}{(l_{k}+r_{k})^{2}}$ (16)
@@
$\mathrm{P}\{|\xi_{k}-\psi[\overline{\mathbf{x}}_{0}]^{k}|\leq(l_{k}+r_{k})\}$ $\geq 1-\frac{\operatorname{var}({\xi_{k}})}{(l_{k}+r_{k})^{2}}$
@@
Since $\psi$ is a row vector with elements summing to one, $\psi[\overline{\mathbf{x}}_{0}]^{k}$ must lie within the range bounded by the maximum and minimum values of $[\overline{\mathbf{x}}_{0}]^{k}$. Consequently, Lemma 4.14 implies that $\xi_{k}$ lies within the interval $[\overline{\mathbf{x}}_{0}]^{k}_{\operatorname{max}}+r_{k}$ to $[\overline{\mathbf{x}}_{0}]^{k}_{\operatorname{min}}-r_{k}$ with a probability of at least $1-\frac{\lambda^{2}/(1-\upsilon^{2})}{(l_{k}+r_{k})^{2}}$.
@@
where $10<\beta<$ 1, $r_{k}$ is a freely chosen parameter for $k=1,2,\ldots,\beta d$, and $\mathcal{A}$ is the convex hull of the initial states of the normal agents. Under PP-ADRC, if the condition in (6) holds, $\overline{\mathcal{G}}(0),\overline{\mathcal{G}}(1),\overline{\mathcal{G}}(2),\ldots$ is repeatedly reachable, and $(1-\beta)$ fraction of the dimensions of the state vector are not added with noises, the probability $\mathrm{P}\{\xi\in\mathcal{C}\}$ that the final value lies within the convex hull $\mathcal{C}$ satisfies that
$\mathrm{P}\{\xi\in\mathcal{C}\}\geq\prod\limits_{k=1}^{\beta d}\left[1-\frac{\lambda^{2}/(1-\upsilon^{2})}{(l_{k}+r_{k})^{2}}\right].$ (19)
Proof 4.20.
@@
First, we assume that $b_{i}$ is a constant and then find the lower bound of $\mathrm{dist}(b,c)$. Apparently, when $c_{i}=b_{i}$, $\mathrm{dist}(b,c)$ is lower bounded, which is $\min\limits_{c_{i}}(\mathrm{dist}(b,c))=\sqrt{\sum\limits_{i=1}^{\kappa}\left[b_{i}^{2}(r_{1}^{2}+\ldots+r_{\beta d}^{2})\right]}$. Then, the problem becomes finding the upper bound of $\min\limits_{c_{i}}(\mathrm{dist}(b,c))$, which is denoted by $\max\limits_{b_{i}}\min\limits_{c_{i}}(\mathrm{dist}(b,c))$. Given that $b_{1}+\ldots+b_{\kappa}=1$ and $b_{1}^{2}+\ldots+b_{\kappa}^{2}\leq(b_{1}+\ldots+b_{\kappa})^{2}$, we can know that the upper bound of $b_{1}^{2}+\ldots+b_{\kappa}^{2}$ is $1$ when $b_{i_{0}}=1,\ i_{0}\in{1,2,\ldots,\kappa}\ \text{ and }\ b_{i}=0,\ i\neq i_{0}$. Therefore, we can obtain $\max\limits_{b_{i}}\min\limits_{c_{i}}(\mathrm{dist}(b,c))=\sqrt{r_{1}^{2}+\ldots+r_{\beta d}^{2}}$. According to the definition, $\max\limits_{b_{i}}\min\limits_{c_{i}}(\mathrm{dist}(b,c))$ is exactly the Hausdorff distance $D_{H}(\mathcal{B},\mathcal{C})$ between $\mathcal{B}$ and $\mathcal{C}$. So we can obtain
@@
$\mathrm{P}\{\xi\in\mathcal{C}\}$ $\geq\prod\limits_{k=1}^{\beta d}\left[1-\frac{\operatorname{var}(\xi_{k})}{(l_{k}+r_{k})^{2}}\right]$
@@ 5 Privacy Analysis
In this section, we analyze the $\rho$-CGP of PP-ADRC and compare it with $(\varepsilon,\delta)$-DP. At last, we make a detailed discussion on the relationship between convergence accuracy and privacy.
5.1 $\rho$-CGP in Resilient Vector Consensus
We first investigate the $\rho$-CGP properties of PP-ADRC. As defined in Definition 8, in the resilient vector consensus, the input is the initial states of normal agents $\overline{\mathbf{x}}_{0}$ and the output is the infinite sequence of noisy states $\mathbf{Y}=\{\overline{\mathbf{y}}(t)\}_{t=0}^{\infty}$.
@@ 5.1 $\rho$-CGP in Resilient Vector Consensus
Under PP-ADRC, if the condition in (6) holds, and $\overline{\mathcal{G}}(0),\overline{\mathcal{G}}(1),\overline{\mathcal{G}}(2),\ldots$ is repeatedly reachable, then we achieve convergence under $\rho$-CGP, where $\rho=\frac{n\upsilon^{2}}{2\lambda^{2}[\upsilon^{2}-(1-\gamma_{l})^{2}]}$.
Proof 5.23.
@@
Then, we show that the outputs $\mathbf{Y}_{k,\overline{\mathbf{x}}_{0}}(\mathbf{N}_{k})$ and $\mathbf{Y}_{k,\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N}_{k}^{\prime})$ share the same distribution under the above definitions. This is proven by the method of mathematical induction. From equation (3), we can derive $y_{i}^{\prime}(0)=x_{i}^{\prime}(0)+\eta_{i}^{\prime}(0)=x_{i}(0)+\delta_{i}+\eta_{i}(0)-\delta_{i}=y_{i}(0),i=1,2,\cdots,\overline{n}$. We can conclude that for every normal agent $i$, the calculated centerpoints $s_{i}(0)$ and $s_{i}^{\prime}(0)$ are the same. According to equation (4), we can derive $x_{i}^{\prime}(1)=x_{i}(1)+(1-\gamma_{i}(0))\delta_{i}$. By induction, we can easily obtain that $x_{i}^{\prime}(h)=x_{i}(h)+\prod_{t=0}^{h-1}\left[1-\gamma_{i}(t)\right]\delta_{i},h=1,2,\cdots,k$, followed by $y_{i}(h)=y_{i}^{\prime}(h)$. It means that $\mathbf{Y}_{k,\overline{\mathbf{x}}_{0}}(\mathbf{N}_{k})=\mathbf{Y}_{k,\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N}_{k}^{\prime})$, indicating that $\mathbf{N}_{k}^{\prime}\in R_{k}^{\prime}$ and we successfully build a bijective correspondence. For any $\mathbf{N}_{k}^{\prime}\in R_{k}^{\prime}$, there exists $(\mathbf{N}_{k},\Delta\mathbf{N}_{k})\in R_{k}\times(\mathbb{R}^{\overline{n}\times d})^{k+1}$ such that $\mathbf{N}_{k}^{\prime}=\mathbf{N}_{k}+\Delta\mathbf{N}_{k}$. Therefore, we derive the integral by changing variables $\mathrm{P}\{\mathbf{N}\in\Omega|\mathbf{Y}_{\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N})\in\mathcal{O}\}=\lim\limits_{k\to\infty}\int_{R_{k}}\mathsf{f}_{\overline{n}(k+1)}(\mathbf{N}_{k}+\Delta\mathbf{N}_{k})\mathrm{d}\mathbf{N}_{k},$ indicating that the distribution $\mathbf{y}_{\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N})$ can be expressed as $\lim\limits_{k\rightarrow\infty}\mathsf{f}_{\overline{n}(k+1)}(\mathbf{N}_{k}+\Delta\mathbf{N}_{k})$.
@@
$=\lim_{k\to\infty}\frac{1}{\alpha-1}\mathrm{log}\bigg{\{}\int_{R_{k,1}}\int_{R_{k,2}}\cdots\int_{R_{k,\overline{n}}}$
$\prod\limits_{h=0}^{k}\prod\limits_{i=1}^{\overline{n}}\bigg{[}\frac{1}{\sqrt{(2\pi\sigma^{2}(h))^{d}}}\mathrm{exp}\left(-\frac{||\eta_{i}(h)-(1-\alpha)\Delta\eta_{i}(h)||^{2}_{2}-||(1-\alpha)\Delta\eta_{i}(h)||^{2}_{2}+(1-\alpha)||\Delta\eta_{i}(h)||^{2}_{2}}{2\sigma^{2}(h)}\right)\bigg{]}$
$\mathrm{d}\eta_{1}(h)\mathrm{d}\eta_{2}(h)\cdots\mathrm{d}\eta_{\overline{n}}(h)\bigg{\}}$
$=\lim_{k\to\infty}\frac{1}{\alpha-1}\mathrm{log}\prod\limits_{h=0}^{k}\{\prod\limits_{i=1}^{\overline{n}}\underset{\eta_{i}(h){\sim}G((1-\alpha)\Delta\eta_{i}(h),\sigma^{2}(h)I_{d})}{\mathbb{E}}\bigg{[}\mathrm{exp}\left(-\frac{-||(1-\alpha)\Delta\eta_{i}(h)||^{2}_{2}+(1-\alpha)||\Delta\eta_{i}(h)||^{2}_{2}}{2\sigma^{2}(h)}\right)\bigg{]}\}$
$=\lim_{k\to\infty}\sum\limits_{h=0}^{k}\bigg{[}\sum\limits_{i=1}^{\overline{n}}\frac{\alpha||\Delta\eta_{i}(h)||^{2}_{2}}{2\lambda^{2}\upsilon^{2h}}\bigg{]}\leq\bigg{[}\frac{\alpha\overline{n}}{2\lambda^{2}}+\lim_{k\to\infty}\sum\limits_{h=1}^{k}\frac{\alpha\overline{n}(1-\gamma_{l})^{2h}}{2\lambda^{2}\upsilon^{2h}}\bigg{]}\mathrm{dist}(\overline{\mathbf{x}}_{0},\overline{\mathbf{x}}_{0}^{\prime})^{2}$
$\leq\alpha\bigg{[}\frac{n\upsilon^{2}}{2\lambda^{2}[\upsilon^{2}-(1-\gamma_{l})^{2}]}\bigg{]}\mathrm{dist}(\overline{\mathbf{x}}_{0},\overline{\mathbf{x}}_{0}^{\prime})^{2}.$
Here, we use $n$ instead of $\overline{n}$, as $n$ is a known parameter. It is worth noting that, as mentioned earlier, the information exchange during each iteration results in decaying leakage of the initial states. As derived from the inequality above, there is a decay factor of $(1-\gamma_{l})$. Therefore, adding a decaying noise is sufficient to achieve privacy preservation.